OpenStudy (anonymous):

What is the solution to the following system using Cramer's Rule? 4x+6y=19 8x-5y=-13 I got two different answers that are not choices!

OpenStudy (turingtest):

what did you get for detA ?

OpenStudy (anonymous):

i think x = -10.16 and then u can solve for y

OpenStudy (turingtest):

x is not -10.16

OpenStudy (anonymous):

so i went det[19,-131;6,-5]/det[4,8;6,-5]

OpenStudy (turingtest):

\[A=\left[\begin{matrix}4 & 6 \\ 8 & -5\end{matrix}\right]\]\[A_1=\left[\begin{matrix}19 & 6 \\ -13 & -5\end{matrix}\right]\]\[A_2=\left[\begin{matrix}4 & 19 \\ 8 & -13\end{matrix}\right]\]\[x=\frac{\det A_1}{\det A}\]\[y=\frac{\det A_2}{\det A}\]take the determinant and use the formula. If you're still having trouble, tell me what you got for each determinant and we will find your mistake.

OpenStudy (anonymous):

is that a -131 or -13

OpenStudy (anonymous):

sorry about that if that is -13 then my answer is wrong